OpenStudy (anonymous):

eight times the sum of a number and 3 is equal to 5

OpenStudy (fibonaccichick666):

Alright, so let's do this a new way.

OpenStudy (fibonaccichick666):

so, are you familiar with he order of operations?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

alright, so you know how if you toss something in parentheses, those items are grouped together and you have to do what is inside first?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

with this, we are going to find the words, sum, difference, product, and/or quotient and put a parenthesis in front of it.

OpenStudy (anonymous):

ok so sum needs a parenthesis

OpenStudy (fibonaccichick666):

then, all of those will be structured like so (sum of _______ and _______) we put a parenthesis after the second item(the one that follows the and

OpenStudy (anonymous):

8(x+3)=5?

OpenStudy (fibonaccichick666):

when we do this, we are giving ourselves a hand for grouping the words where they need to be

OpenStudy (fibonaccichick666):

that is correct

OpenStudy (fibonaccichick666):

try another

OpenStudy (anonymous):

the difference of a number times 4 and 9 equals 7

OpenStudy (fibonaccichick666):

so group first, then write

OpenStudy (anonymous):

x(4-9)=7

OpenStudy (fibonaccichick666):

no, what word comes first?

OpenStudy (anonymous):

the difference

OpenStudy (anonymous):

x-4?

OpenStudy (fibonaccichick666):

then make blanks so actually underline the two things that are being subtracted

OpenStudy (anonymous):

tht would be x and 4

OpenStudy (anonymous):

right

OpenStudy (fibonaccichick666):

the (difference of \(\underline{ \text{a number times 4}}\) and \(\underline{ \text{9 }}\)equals 7

OpenStudy (anonymous):

ahhhhh, x-9

OpenStudy (fibonaccichick666):

no, you have to do the whole first underline

OpenStudy (anonymous):

4x-9

OpenStudy (fibonaccichick666):

there we go

OpenStudy (anonymous):

ok, there are no more questions online

OpenStudy (anonymous):

but you rock thank you so much

OpenStudy (fibonaccichick666):

uh, I'm sure there is a generator somewhere. Hmm

OpenStudy (fibonaccichick666):

A number increased by nine is fifteen

OpenStudy (anonymous):

ok, i'll leave the thread open if you wanna find me more problems, i gotta move on though, i got one more lesson to do before the end of the night

OpenStudy (fibonaccichick666):

Twice a number, decreased by twenty-nine, is seven.

OpenStudy (anonymous):

x*9=15

OpenStudy (fibonaccichick666):

increased doesn't mean mult

OpenStudy (fibonaccichick666):

The quotient of fifty and five more than a number is ten

OpenStudy (anonymous):

x+9?

OpenStudy (fibonaccichick666):

yea

OpenStudy (anonymous):

cool

OpenStudy (fibonaccichick666):

and here is the site I am getting them from. Take a look, http://www.mathgoodies.com/lessons/vol7/equations.html

OpenStudy (fibonaccichick666):

best of luck

OpenStudy (anonymous):

ok, thank you

OpenStudy (fibonaccichick666):

np :)